Institut for Matematik og Datalogi Odense Universitet April 11, 1996

JFB

DM11 – F96 – Lecture 10

Lecture, April 10

We covered sections 5.1 and 5.2 and most of Appendix 3.

Lecture, April 17

We will finish Appendix 3, cover section 5.3, and begin chapter 6.

Lecture, April 24

We will continue in chapter 6.

Announcement 1

On Wednesday, April 24, the lecture will be held in U37, instead of U110.

Announcement 1

On Thursday, April 25, the discussion section D2 from U13 will be held in U22.

Problems to be discussed in week 17

• Solve the recurrence from problem 17 of section 5.1.

• 18, 19, 20, and 21 from section 5.2.

• 3, 6, 9, 10, and 12 from Appendix 3.

• 14, 16, 17, 18, 19, 20, 21 from section 5.3.

Hint: use induction for 17 and 19. In 19, there is typo. The symbols C_{2} and c_{2} should be the same.

• Prove that for |ax| < 1, 1/(1 − ax)^{m+1} = ∑_{k=0}^{∞} ( \begin{matrix}m + k \\ m \end{matrix} ) a^{k}x^{k}.